News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 26 October 2001 10.30am EDT

Last week the observing schedule was halted twice due to high radiation associated with solar flare activity. The loads were halted on Oct 21 at 11:38am EDT, resumed on Oct 23 at 4:00am EDT and halted a second time on Oct 25 at 7:17pm. In both cases the loads were halted through a ground command to execute the SI Safing SCS 107 since the projected orbital radiation dose was significantly higher than the threshold allowed for ACIS. The loads are being replanned to resume on Oct 27 at 6:40am EDT. Observations of Puppis N Knot, PSS 0902+1640 and PSR B0833-45 were impacted by the first event, and observations of MG J0414+0534, HS0218+3707, PSS0134+3307, 2MASSJ222202+1952, M17 and SAX J1747.0-285 were impacted by the second event. These targets will be rescheduled in a future load.

A Chandra press release was issued on 18 Oct describing observations of the Herbig-Haro object HH2 in the Orion Nebula. The results show this type of object does produce X-rays and proves a theory of how the objects form. For the details see: http://chandra.harvard.edu/press/01_releases/press_101901.html.

The schedule of observations starting from the re-plan intercept through the coming week is shown below.

------------------------------------------
SN1998BW                ACIS-S       Oct 27
Radiation Belts
BHC Transient           ACIS-S/HETG  Oct 28
HD 93497                ACIS-S/HETG  Oct 29
RR Pic                  ACIS-S
IRAS 05189-2524         ACIS-S       Oct 30
Radiation Belts
IRAS 20551-4250         ACIS-S       Oct 31
LBQS 0025-0151          ACIS-S
M31                     HRC-I
53W002 Cluster          ACIS-S       Nov 1 
Markarian 266           ACIS-S
UGC 4305                ACIS-S       Nov 2
Radiation Belts
S106 FIR                ACIS-I
NGC 3665 Group          ACIS-I       Nov 3
ABELL 611               ACIS-S
ZWICKY 2701             ACIS-S       Nov 4
Puppis Bright E Knot    ACIS-S
PSR0628-28              ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports